Mauro Products distributes a single product, a woven basket whose selling price is $18 per unit and whose variable expense is $14 per unit. The company’s monthly fixed expense is $10,400.

Required:

1. Calculate the company’s break-even point in unit sales.

2. Calculate the company’s break-even point in dollar sales. (Do not round intermediate calculations.)

3. If the company's fixed expenses increase by $600, what would become the new break-even point in unit sales? In dollar sales? (Do not round intermediate calculations.)
Answer #1

1. Break even point = Fixed cost/(sales - variable cost)

= 10400/(18-14)

= 2600 units

2. Break even point = 2600*18 = $46800

3. Break even point = (10400+600)/(18-14) = 2750 units

Break even point = 2750*18 = $49500